OpenStudy (anonymous):

Again help plzz! :D 4x-2 (x-5)=18

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

4x-2x+10=18 2x=18-10 2x=8 x=4

OpenStudy (anonymous):

Thankss soo muchh

OpenStudy (anonymous):

yw